# Properties of observable mixed inertial/gravito-inertial modes in $\gamma$ Doradus stars

### v1

@@ 4.3 Relative errors made on the estimation of $N_{0}$ and $s_{*}$
The relative error of $s_{*}$ is far smaller and does not get greater than $\sim 0.7\%$. It tends to decrease as the stars evolves. As dips are not currently observed in stars as evolved as the 3t model, we also estimated the parameters $A$ and $\beta$ using only the 1z and 2m models (see the results in Table 5). The relative error made on $N_{0}$ and $s_{*}$ for the series ($\ell=1$, $m=-1$) using these new estimates of $A$ and $\beta$ are shown in Fig. 16. fl The relative error of $N_{0}$ is now always inferior to 6% and the relative error on $s_{*}$ is inferior to 0.2% except for the case 1z and $\Omega/\Omega_{\mathrm{K}}=0.5$, where it reaches $\sim$ 1%. The same approach is used for the series ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$) and reported in Appendix C.
Figure 15: Relative errors for the series ($\ell=1$, $m=-1$) on the estimated $N_{0}^{\mathrm{calc}}$ (upper panel) and $s_{*}^{\mathrm{calc}}$ (lower panel) using the three models 1z, 2m and 3t.
Table 5: Values of $A$ and $\beta$ obtained by fitting the relations $(s_{c}/s_{*}-1)N_{0}=A\Omega$ and $\sigma N_{0}=\beta\Omega$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$), by only taking into account the models 1z and 2m together.
@@
We showed that, for observable dips, our empirical method can predict accurately the frequency of the resonant inertial mode in the core and provides a reasonable estimate of the jump in the Brunt-Väisälä frequency at the interface between the core and the envelope. Both of them are correlated to the evolutionary stage of the star. In particular, the frequency of the resonant inertial mode is only sensitive to the structure of the core. It is thus independent of mixing and diffusion processes occurring in the overlying radiative layers, which, in contrast, may deeply affect the shape of the Brunt-Väisälä frequency profile.
Further investigations are needed in order to characterise how the frequencies of inertial modes are depending on the characteristics of the convective core. Inertial modes could thus be a powerful tool to obtain constraints on convective cores and so precise estimations of stellar ages.
@@ 5 Approximate analytical models
We present the dip model of Tokuno & Takata (2022) and a new improved version of this model and compare them with the properties of the dip obtained by numerical calculations in the previous sections.
@@ 5.1 Model of Tokuno & Takata (2022)
Tokuno & Takata (2022) proposed an analytical model of the dips in $\Delta P=f(s)$ provoked by the coupling between gravito-inertial and inertial oscillations. It is based on three main assumptions: the density is uniform in the convective core, the gravito-inertial oscillations are described by the TAR and the ratio between the (uniform) rotation rate and the Brunt-Väisälä frequency at the bottom of the radiative zone is small. The assumption of uniform density is obviously unrealistic but it enables to get analytical solutions of the inertial oscillations in the core. In the absence of coupling with the radiative envelope, inertial modes in the uniform density core are identified by their spin parameter, $s_{*}$, and the degree and azimuthal order, $(\ell_{i},m)$, of the Legendre polynomial involved in the solution (see Appendix A). Under the TAR, the latitudinal part of gravito-inertial $(\ell,m)$ modes is characterised by the Hough function $H_{k}^{m}(\theta,s)$ associated with the eigenvalue $\Lambda_{k}^{m}(s)$, where $k=\ell-|m|$ (see Eq. 33). The model results depend on the continuity of the density and the Brunt-Väisälä frequency at the interface between the convective core and the radiative zone. For all stellar models of $\gamma$ Dor stars presented in Sect. 2.1, the density is continuous at the convective/radiative interface while the Brunt-Väisälä is discontinuous, the jump from zero on the convective side to $N_{0}$ on the radiative side increasing with age. According to Tokuno & Takata (2022), the coupling between the $(\ell,m)$ series of gravito-inertial modes and a $(\ell_{i},m)$ inertial mode produces a Lorentzian-shaped dip in the $\Delta P=f(s)$ that depends on two parameters: its width [^2] The expressions of $\sigma$ and $s_{c}$ in the case considered here are not explicitly written in Tokuno & Takata (2022), but can easily be derived from their equations (89), (90) and (91) with $\Delta\rho=0$ and replacing $\tilde{F}$ by $F$ given by their equation (58).,
@@
A first success of this model is that the Lorentzian form of the dips is fully consistent with our numerical results (see Sect. 3.1).
@@
The expressions of $\sigma$ and $s_{c}$ indicate that the dip parameters depend on the ratio $\Omega/N_{0}$, on the gravito-inertial oscillations through $\Lambda_{k}^{m}$, and on the inertial oscillation through $s_{*}$ and $R(s_{*})=\frac{2P_{\ell_{i}}^{m}(1/s_{*})}{{\gamma^{\prime}}_{\ell_{i}}^{m}(s_{*})}$. Tokuno & Takata (2022) argued that these expressions remain relevant in the case of a realistic core with a non-uniform density, although the quantities related to the inertial oscillation that is $s_{*}$ and the function $R(s)$ are no longer known analytically. We can nevertheless determine $s_{*}$ numerically by computing the inertial modes of stellar models truncated at the boundary of the convective core where the radial displacements vanish (see Appendix B). This allows us to test the $s_{c}=s_{*}$ prediction against the numerical determination of $s_{c}$ for the dips studied in this paper. As shown in Fig. 10, our numerical results are not consistent with $s_{c}=s_{*}$.
@@
The theoretical expression of $\sigma$ can not be fully tested with our numerical results because the function $R(s)$, which is derived from the analytical form of free inertial oscillations at the convective/radiative interface, is unknown for cores of variable density. Nevertheless, the proportionality $\sigma\propto\Omega$ can be tested by considering the evolution of dip widths with rotation for a given star model and $(\ell,m)$ series. Figure 12 shows that $\sigma\propto\Omega$ is indeed in agreement with our numerical results despite the small deviations observed in the case of the most evolved star model. Tokuno & Takata (2022) also treated the case of a continuous Brunt-Väisälä frequency profile at the core interface and found a different scaling of the dip width with the rotation. We note that even the small jump of the Brunt-Väisälä in the 1z model produces an evolution of $\sigma$ corresponding to the discontinuous case. A questionable property of the theoretical $\sigma$ is that each inertial mode having the same $m$ and equatorial parity than the gravito-inertial modes of a $(\ell,m)$ series produces a dip of non negligible width. As already discussed in Tokuno & Takata (2022), this property is at odds with Ouazzani et al. (2020) phenomenological model where an efficient resonance requires a non-negligible geometrical matching between the gravito-inertial and inertial modes at the convective/radiative interface. This matching selects the significantly resonant inertial modes and thus the dips in the $\Delta P=f(s)$, a prediction that was successfully tested with full numerical calculations by Ouazzani et al. (2020).
5.2 An improved model
@@
In this subsection, we present a new improved analytical model of the dips studied in this paper. This model is similar to Tokuno & Takata (2022) model in that it uses the same three basic assumptions: uniform-density core, TAR gravito-inertial oscillations and $\Omega/N_{0}\ll 1$, but it goes a step further by improving the geometrical matching of the inertial and gravito-inertial oscillations at the convective/radiative interface. This matching is indeed required by the continuity condition on the radial displacement and the pressure perturbations at this interface. Figure 17 shows the latitudinal profiles of the gravito-inertial mode and the inertial mode involved in the $(\ell=1,m=-1)$ vs $(\ell_{i}=3,m=-1)$ resonance. It is clear that the continuity condition cannot be realised by only considering these two modes as in Tokuno & Takata (2022). We improve the matching by allowing more complex eigenfunctions on both side of the interface. The detail of the derivation is given in Appendix D.
@@
Figure 17: Latitudinal profiles of the $\ell_{i}=3,m=-1,s_{*}=11.32$ inertial mode ($P_{3}^{-1}(\cos\theta)$, blue) and $\ell=1,m=-1$ gravito-inertial modes near $s_{*}$ ($H_{0}^{-1}(\theta,s_{*})$, orange) at the convective/radiative interface. Mixed inertial/gravito-inertial modes continuous at the interface can hardly be constructed with these two modes only.
@@
We find that the coupling between the $(\ell=1,m=-1),(\ell=2,m=-2),(\ell=1,m=0)$ gravito-inertial modes and respectively the $(\ell_{i}=3,m=-1),(\ell_{i}=4,m=-2),(\ell_{i}=3,m=0)$ inertial modes produces Lorentzian-shaped dips in the $\Delta P=f(s)$ relation whose width and central spin parameter are:
@@
$f_{k,\ell_{i}}^{m}(s_{*})=-\frac{c_{k,\ell_{i}}^{m}(s_{*})c_{k-2,\ell_{i}-2}^{m}(s_{*})}{c_{k,\ell_{i}-2}^{m}(s_{*})c_{k-2,\ell_{i}}^{m}(s_{*})},$ (18)
where $c^{m}_{k,\ell_{i}}$ characterises the geometrical coupling between inertial and gravito-inertial oscillations along the convective/radiative interface:
@@
$\tilde{P}_{\ell_{i}}^{m}(\cos\theta)$ and $H_{k}^{m}(\theta,s)$ being normalised Legendre polynomial and Hough functions. In addition to $P_{\ell_{i}}^{m}$ and $H_{k}^{m}$, these mixed gravito-inertial/inertial modes involve $P_{\ell_{i}-2}^{m}$ and $H_{k-2}^{m}$. In Appendix D, we argue this is a good approximation for the first dip of the $(\ell=1,m=-1)$, $(\ell=2,m=-2)$ and $(\ell=1,m=0)$ series studied in this paper. This model cannot be generalised to all dips as we expect some resonances to produce more complex mixed modes.
@@
The dip properties predicted by this model have important differences with Tokuno & Takata (2022) and these differences appear to improve the model. First, the width $\sigma$ is multiplied by the factor $f_{k,\ell_{i}}^{m}/(1+f_{k,\ell_{i}}^{m})$. As $f_{k,\ell_{i}}^{m}$ is proportional to the coupling coefficient $c_{k,\ell_{i}}(s_{*})$ introduced in Ouazzani et al. (2020), this new expression of $\sigma$ can account for the fact that significant dips only occur when the resonance involve modes with a non-negligible geometrical matching at the convective/radiative interface. Second, the new form of the central spin parameter is now consistent with the numerical results as $s_{c}-s_{*}\propto\Omega$.
@@
When extrapolated to a realistic core with a non-uniform density, our model suggests that the first dip of $(\ell=1,m=-1),(\ell=2,m=-2),(\ell=1,m=0)$ $(\ell,m)$ series can be expressed as:
$\sigma=$ $\beta\frac{\Omega}{N_{0}}$ (20)
$s_{c}=$ $s_{*}+\alpha\frac{\Omega}{N_{0}}$ (21)
@@
where $\beta$, $s_{*}$ and $\alpha$ depend on both the gravito-inertial modes and on the resonant inertial mode. For a fixed $(\ell,m)$, these parameters only depend on the properties of the inertial mode that is on the radial gradient of the density profile in the convective core (see Eq. (1) in Ouazzani et al., 2020). In the absence of analytical expressions, they can be derived by fitting the dips computed for different star models. In Sect. 4, using three star models, it was found that $\beta$ remain roughly constant while $\alpha$ is approximately proportional to $1/s_{*}$. An open question for future more detailed empirical relations is whether $\beta$ and $\alpha$ can indeed be described by functions of $s_{*}$ only. In that case, investigating the link between the seismic observable $s_{*}$ and the star properties (age, mass, overshooting,…) would be much facilitated. Computing $s_{*}$ for given stellar model is indeed much easier than computing the frequency pattern and the dip parameters of a $(\ell,m)$ mode series.
@@
Because of the uniform density assumption, a quantitative agreement between the theoretical $\sigma$ and $s_{c}$ and the numerical ones is not expected. We nevertheless compared them by calculating the theoretical values of $\beta_{\rm th}(s_{*})$, $\alpha_{\rm th}(s_{*})$ and $A_{\rm th}=\alpha_{\rm th}(s_{*})s_{*}$. Table 6 presents these parameters for the three dips studied. It shows that, despite the uniform density assumption, the theoretical values are in the range of those obtained by numerical calculations (see Tables 2, 3, 4 and 5). This is consistent with the fact that the spin parameter and the surface distribution of the inertial modes are not strongly modified by the core density stratification (see Appendix B). We also note that, as expected, for $\alpha$ the deviations from the uniform-density values increase for more evolved star models with higher density contrasts.

### v2

@@ 4.3 Relative errors made on the estimations of $N_{0}$ and $s_{*}$
The relative error of $s_{*}$ is far smaller and does not become greater than $\sim 0.7\%$. It tends to decrease as the stars evolves. As dips are not currently observed in stars as evolved as the 3t model, we also estimated the parameters $A$ and $\beta$ using only the 1z and 2m models (see the results in Table 5). The relative error made on $N_{0}$ and $s_{*}$ for the series ($\ell=1$, $m=-1$) using these new estimates of $A$ and $\beta$ are shown in Fig. 16. The relative error of $N_{0}$ is always inferior to 6%, and the relative error on $s_{*}$ is inferior to 0.2%, except for the case of 1z and $\Omega/\Omega_{\mathrm{K}}=0.5$, where it reaches $\sim$ 1%. The same approach was used for the series ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$), and the results are reported in Appendix C.
Figure 15: Relative errors for the series ($\ell=1$, $m=-1$) on the estimated $N_{0}^{\mathrm{calc}}$ (upper panel) and $s_{*}^{\mathrm{calc}}$ (lower panel) using the three models 1z, 2m, and 3t.
Table 5: Values of $A$ and $\beta$ when only taking into account the models 1z and 2m together.
@@
 [^5] Values of $A$ and $\beta$ are obtained by fitting the relations $(s_{c}/s_{*}-1)N_{0}=A\Omega$ and $\sigma N_{0}=\beta\Omega$ for the series of modes ($\ell=1$, $m=-1$), ($\ell=1$, $m=0$), and ($\ell=2$, $m=-2$).
@@
We showed that for observable dips, our empirical method can accurately predict the frequency of the resonant inertial mode in the core, and it provides a reasonable estimate of the jump in the Brunt-Väisälä frequency at the interface between the core and the envelope. Both $N_{0}$ and $s_{*}$ are correlated to the evolutionary stage of the star. In particular, the frequency of the resonant inertial mode is only sensitive to the structure of the core. It is thus independent of the mixing and diffusion processes occurring in the overlying radiative layers, which, in contrast, may deeply affect the shape of the Brunt-Väisälä frequency profile.
Further investigations are needed in order to characterise how the frequencies of inertial modes depend on the characteristics of the convective core. Inertial modes could be a powerful tool in obtaining constraints on convective cores and thus precise determinations of stellar ages.
@@ 5 Approximate analytical models
We present the dip model of Tokuno & Takata (2022) as well as a newly improved version of the model and compare them with the properties of the dip obtained by numerical calculations in the previous sections.
@@ 5.1 Model of Tokuno & Takata (2022)
Tokuno & Takata (2022) proposed an analytical model of the dips in $\Delta P=f(s)$ provoked by the coupling between gravito-inertial and inertial oscillations. It is based on three main assumptions: the density is uniform in the convective core, the gravito-inertial oscillations are described by the TAR, and the ratio between the (uniform) rotation rate and the Brunt-Väisälä frequency at the bottom of the radiative zone is small. The assumption of uniform density is obviously unrealistic, but it enables one to get analytical solutions of the inertial oscillations in the core. In the absence of coupling with the radiative envelope, inertial modes in the uniform density core are identified by their spin parameter, $s_{*}$, and the degree and azimuthal order, $(\ell_{i},m)$, of the Legendre polynomial involved in the solution (see Appendix A). Under the TAR, the latitudinal part of gravito-inertial $(\ell,m)$ modes is characterised by the Hough function $H_{k}^{m}(\theta,s)$ associated with the eigenvalue $\Lambda_{k}^{m}(s)$, where $k=\ell-|m|$ (see Eq. 33). The model results depend on the continuity of the density and the Brunt-Väisälä frequency at the interface between the convective core and the radiative zone. For all stellar models of $\gamma$ Dor stars presented in Sect. 2.1, the density is continuous at the convective-radiative interface, while the Brunt-Väisälä is discontinuous. Indeed, the jump from zero on the convective side to $N_{0}$ on the radiative side increases with age. According to Tokuno & Takata (2022), the coupling between the $(\ell,m)$ series of the gravito-inertial modes and a $(\ell_{i},m)$ inertial mode produces a Lorentzian-shaped dip in the $\Delta P=f(s)$ that depends on two parameters, namely its width, [^6] The expressions of $\sigma$ and $s_{c}$ in the case considered here are not explicitly written in Tokuno & Takata (2022) but can easily be derived from their equations (89), (90), and (91) with $\Delta\rho=0$ and replacing $\tilde{F}$ by $F$ given by their equation (58).
@@
A first success of this model is that the Lorentzian form of the dips is fully consistent with our numerical results (see Sect. 3.1). The expressions of $\sigma$ and $s_{c}$ indicate that the dip parameters depend on the ratio $\Omega/N_{0}$, on the gravito-inertial oscillations through $\Lambda_{k}^{m}$, and on the inertial oscillation through $s_{*}$ and $R(s_{*})=2P_{\ell_{i}}^{m}(1/s_{*})/{\gamma^{\prime}}_{\ell_{i}}^{m}(s_{*})$. Tokuno & Takata (2022) argued that these expressions remain relevant in the case of a realistic core with a non-uniform density, although the quantities related to the inertial oscillation, that is $s_{*}$ and the function $R(s)$, are no longer known analytically. We can nevertheless determine $s_{*}$ numerically by computing the inertial modes of stellar models truncated at the boundary of the convective core where the radial displacements vanish (see Appendix B). This allowed us to test the $s_{c}=s_{*}$ prediction against the numerical determination of $s_{c}$ for the dips studied in this paper. As shown in Fig. 10, our numerical results are not consistent with $s_{c}=s_{*}$.
@@
The theoretical expression of $\sigma$ cannot be fully tested with our numerical results because the function $R(s)$, which is derived from the analytical form of free inertial oscillations at the convective-radiative interface, is unknown for cores of variable density. Nevertheless, the proportionality $\sigma\propto\Omega$ can be tested by considering the evolution of dip widths with rotation for a given star model and $(\ell,m)$ series. Figure 12 shows that $\sigma\propto\Omega$ is indeed in agreement with our numerical results despite the small deviations observed in the case of the most evolved star model. Tokuno & Takata (2022) also treated the case of a continuous Brunt-Väisälä frequency profile at the core interface and found a different scaling of the dip width with the rotation. We note that even the small jump of the Brunt-Väisälä in the 1z model produces an evolution of $\sigma$ corresponding to the discontinuous case. A questionable property of the theoretical $\sigma$ is that each inertial mode with the same $m$ and equatorial parity as the gravito-inertial modes of a $(\ell,m)$ series produces a dip of non-negligible width. As already discussed in Tokuno & Takata (2022), this property is at odds with the Ouazzani et al. (2020) phenomenological model, where an efficient resonance requires a non-negligible geometrical matching between the gravito-inertial and inertial modes at the convective-radiative interface. This matching selects the significantly resonant inertial modes and thus the dips in the $\Delta P=f(s)$, a prediction that was successfully tested with full numerical calculations by Ouazzani et al. (2020).
5.2 The improved model
@@
In this subsection, we present a newly improved analytical model of the dips studied in this paper. This model is similar to the Tokuno & Takata (2022) model in that it uses the same three basic assumptions: uniform density core, TAR gravito-inertial oscillations, and $\Omega/N_{0}\ll 1$. However, it goes a step further by improving on the geometrical matching of the inertial and gravito-inertial oscillations at the convective-radiative interface. This matching is indeed required by the continuity condition on the radial displacement and the pressure perturbations at this interface. Figure 17 shows the latitudinal profiles of the gravito-inertial mode and the inertial mode involved in the $(\ell=1,m=-1)$ versus $(\ell_{i}=3,m=-1)$ resonance. It is clear that the continuity condition cannot be realised by only considering these two modes, as is done in Tokuno & Takata (2022). We improved the matching by allowing more complex eigenfunctions on both sides of the interface. The detail of the derivation is given in Appendix D.
@@
Figure 17: Latitudinal profiles of the $\ell_{i}=3,m=-1,s_{*}=11.32$ inertial mode ($P_{3}^{-1}(\cos\theta)$, blue) and $\ell=1,m=-1$ gravito-inertial modes near $s_{*}$ ($H_{0}^{-1}(\theta,s_{*})$, orange) at the convective-radiative interface. Mixed modes that are continuous at the interface can hardly be constructed with these two modes only.
@@
We find that the coupling between the $(\ell=1,m=-1),(\ell=2,m=-2),(\ell=1,m=0)$ gravito-inertial modes and respectively the $(\ell_{i}=3,m=-1),(\ell_{i}=4,m=-2),(\ell_{i}=3,m=0)$ inertial modes produces Lorentzian-shaped dips in the $\Delta P=f(s)$ relation whose widths and central spin parameters are described by
@@
$f_{k,\ell_{i}}^{m}(s_{*})=-\frac{c_{k,\ell_{i}}^{m}(s_{*})c_{k-2,\ell_{i}-2}^{m}(s_{*})}{c_{k,\ell_{i}-2}^{m}(s_{*})c_{k-2,\ell_{i}}^{m}(s_{*})}.$ (18)
Here, $c^{m}_{k,\ell_{i}}$ characterises the geometrical coupling between inertial and gravito-inertial oscillations along the convective-radiative interface:
@@
with $\tilde{P}_{\ell_{i}}^{m}(\cos\theta)$ and $H_{k}^{m}(\theta,s)$ being the normalised Legendre polynomial and Hough functions. In addition to $P_{\ell_{i}}^{m}$ and $H_{k}^{m}$, these mixed modes involve $P_{\ell_{i}-2}^{m}$ and $H_{k-2}^{m}$. In Appendix D, we argue this is a good approximation for the first dip of the $(\ell=1,m=-1)$, $(\ell=2,m=-2)$, and $(\ell=1,m=0)$ series studied in this paper. This model cannot be generalised to all dips, as we expect some resonances to produce more complex mixed modes.
@@
The dip properties predicted by this model have important differences with Tokuno & Takata (2022), and these differences appear to improve the model. First, the width $\sigma$ is multiplied by the factor $f_{k,\ell_{i}}^{m}/(1+f_{k,\ell_{i}}^{m})$. As $f_{k,\ell_{i}}^{m}$ is proportional to the coupling coefficient $c_{k,\ell_{i}}(s_{*})$ introduced in Ouazzani et al. (2020), this new expression of $\sigma$ can account for the fact that significant dips only occur when the resonance involves modes with a non-negligible geometrical matching at the convective-radiative interface. Second, the new form of the central spin parameter is now consistent with the numerical results as $s_{c}-s_{*}\propto\Omega$.
@@
When extrapolated to a realistic core with a non-uniform density, our model suggests that the first dip of the $(\ell=1,m=-1)$, $(\ell=2,m=-2)$, and $(\ell=1,m=0)$ series can be expressed as
$\sigma=$ $\beta\frac{\Omega}{N_{0}},$ (20)
$s_{c}=$ $s_{*}+\alpha\frac{\Omega}{N_{0}},$ (21)
@@
where $\beta$, $s_{*}$, and $\alpha$ depend on both the gravito-inertial modes and on the resonant inertial mode. For a fixed $(\ell,m)$, these parameters only depend on the properties of the inertial mode, that is, on the radial gradient of the density profile in the convective core (see Eq. (1) in Ouazzani et al., 2020). In the absence of analytical expressions, they can be derived by fitting the dips computed for different star models. In Sect. 4, using three star models, it was found that $\beta$ remains roughly constant, while $\alpha$ is approximately proportional to $1/s_{*}$. An open question for future and more detailed empirical relations is whether $\beta$ and $\alpha$ can indeed be described by functions of $s_{*}$ only. In that is the case, investigating the link between the seismic observable $s_{*}$ and the star properties (age, mass, overshooting, etc.) would be much facilitated. Computing $s_{*}$ for a given stellar model is indeed much easier than computing the frequency pattern and the dip parameters of a $(\ell,m)$ mode series.
@@
Because of the uniform density assumption, a quantitative agreement between the theoretical and numerical values of $\sigma$ and $s_{c}$ is not expected. We nevertheless compared them by calculating the theoretical values of $\beta_{\rm th}(s_{*})$, $\alpha_{\rm th}(s_{*})$, and $A_{\rm th}=\alpha_{\rm th}(s_{*})s_{*}$. Table 6 presents these parameters for the three dips we studied. It shows that despite the uniform density assumption, the theoretical values are in the range of those obtained by numerical calculations (see Tables 2, 3, 4, and 5). This is consistent with the fact that the spin parameter and the surface distribution of the inertial modes are not strongly modified by the core density stratification (see Appendix B). We also note that as expected for $\alpha$, the deviations from the uniform density values increase for more evolved star models with higher density contrasts.